# Mergers of compact objects with cores of massive stars: evolution, r-process and multi-messenger signatures

### v1

@@ 4 R-process nucleosynthesis and enrichment
The possible formation site(s) of r-process elements remain a highly debated subject (e.g., Hotokezaka et al. 2018; Siegel et al. 2019; Ji et al. 2019; Banerjee et al. 2020; Chen et al. 2024a; Chen et al. 2024b; Tsujimoto 2024). In 2017, LIGO detected the first GW signature from binary NS (BNS) mergers, commonly known as GW170817 (Abbott et al., 2017a). The electromagnetic counterpart that was associated with this GWs event (kilonova AT2017gfo; Abbott et al. 2017b) showed traces of lanthanides in their equatorial outflow (Kasen et al., 2017), hence confirming r-process nucleosynthesis in BNS mergers. Despite this groundbreaking progress, which was further strengthen by an additional discovery of a kilonova that might be associated with another BNS merger with JWST (Levan et al., 2024), it is not clear whether BNS mergers are the only site of heavy elements formation in the Universe. The long delay times until the merger and subsequent r-process nucleosynthesis, reaching tens of Gyrs (e.g., Belczynski et al. 2018), might imply there are not enough mergers to account for heavy elements formation in the early Universe (e.g., Côté et al. 2019; Kobayashi et al. 2023. See Maoz & Nakar 2024 and Beniamini & Piran 2024 for a different view). Moreover, these long delay times combined with the two natal-kicks that the binary system receives following each SN explosion (see left panel of Fig. 1) might drive the BNS outside of ultra-faint dwarf (UFD) galaxies before they manage to merge, implying that they are not able to enrich them with heavy elements (e.g., Bonetti et al. 2019). However, recent observations of the UFD galaxy Reticulum II show r-process enhanced gas (e.g., Ji et al. 2016; Roederer et al. 2016; Ji et al. 2023; Simon et al. 2023), highlighting the need for additional r-process nucleosynthesis sites (see Beniamini et al. 2016 for a another view.)
@@
Papish et al. (2015) suggest that r-process nucleosynthesis can occur during the merger of an NS with the core of a giant star. As the NS spirals through the core of the giant, it might accrete mass at very high rates (see section 3 for a discussion about the accretion rate). The vast accretion destroys the core forming a thick accretion disk around the NS that can self neutronize through electron captures due to the extremely high temperatures and densities in the disk, providing a neutron-rich environment where heavy elements can be formed via rapid neutron captures by isotopes that come from outer cooler layers of the disk. Grichener & Soker (2019) showed that under their model assumptions the matter of the accretion disk could be neutron-rich enough for r-process nucleosynthesis if the merger occurs after helium core depletion, when the core is CO-rich. They perform one-dimensional stellar evolution simulations of several massive stars that engulf an NS in different stages of their evolution and mimic the effect of the NS on the giant by removing envelope mass from the location of the NS outwards. Using the Bondi-Hoyle-Lyttleton approximation (see equation 2), they compute the accretion rate of the NS as it mergers with the core of the giant, and then apply it to estimate the typical temperatures and densities of the material in the disk within the analytical framework of a neutrino cooled accretion disk from Chevalier (1996)
@@
Since the orbital shrinkage leading to mergers of NSs with giants’ cores is driven by dynamical friction during CEE rather than GWs irradiation, they occur on a dynamical timescale and have much shorter delay time from the formation of the binary compared to BNS mergers ( tens of Myrs; from the population synthesis data of Grichener 2023; comparable to the lifetime of the NS progenitor). These shorter delay times offer an advantage in explaining the presence of heavy element in the early Universe. Moreover, NS-core mergers suffer from a single natal-kick induced by the SN that formed the NS, increasing the possibility that they will be retained and produce heavy elements in UFD galaxies (Grichener & Soker, 2022). Grichener (2023) computes the event rates of NS-core mergers with CO-rich cores and find it to be around 1 event per 1000 CCSNe for fiducial model parameters, which is comparable to the rate of BNS mergers inferred by LIGO (Abbott et al., 2023), and sufficient to account for a substantial fraction of observed r-process abundances. Grichener et al. (2022) use a Galactic Chemical Evolution model of the Milky-Way (Kobayashi et al., 2020) and include the contribution of NS-core mergers to r-process nucleosynthesis. They study the time evolution of europium in the Galaxy and find that NS-core mergers (orange dots in Fig. 3) could account for both the average of the scatter at earlier times and the ’evolution knee’ that follows, while BNS mergers (yellow dots in Fig. 3) struggle to do so.
@@ 5.1.1 Merger-driven lightcurves properties
While three-dimensional hydrodynamical simulations that follow the the outflows and outgoing radiation during the merger of the compact object with the core are scarce, simulating the final CEE stages of a compact object and a giant star has value in building intuition. Schrøder et al. (2020) model the in-spiral of a BH inside a giant and calculate the lightcurves expected prior to the merger of the BH with the core accounting for the effects of CSM interaction with a spherically generated shock instead of jets. They continue the simulations until the orbital separation of the binary is about , assuming the subsequent merger is expected to occur rapidly on a timescale of less then a year. They then analyze the ejection of matter during the final stage of CEE and find the resultant CSM distribution . We note that by not modelling the accretion onto the BH, their resultant CSM might be more massive than expected is such a merger. For computational simplicity, they derive a spherically symmetric density distribution of the ejecta and model the CSM interactions using a one-dimensional radiative hydrodynamics code.
@@
As can be seen in Fig 4, Schrøder et al. (2020) find that at early times the lightcurves are mostly dominated by CSM interactions. As in other models of dense CSM close to the giant star, they find fast rising high luminosities that reach plateau over a period of hundreds of days. 333In the case of an extended CSM the decay of the lightcurve can happen over a period of thousands of days (e.g., Smith 2017). Typical models have higher luminosities than type IIp SNe, but still lower than in the case of super-luminous SNe. The top panel of Fig. 4 shows that more energetic mergers result in shorter, more luminous transients with less prominent effect of CSM interactions on the lightcurve as the larger photospheres of the explosive transients are further out from the CSM edge. The lightcurve of the most energetic transient ( ergs) reassembles the shape of a type IIp SN (Schrøder et al., 2020). Higher giant masses (middle panel of Fig. 4) result in lower ejecta velocities that lead to longer transients. Different giant radii (bottom panel of Fig. 4) affect the transients mainly by forming CSMs of different extents. Smaller giants lead to more compact CSMs that can affect the lightcurve through interactions at earlier times. Such explosive transients could be detected by current facilities, such as ZTF (Bellm et al., 2019) and ASAS-SN (Kochanek et al., 2017), as well as with future programs like LSST. The CSM interactions involved in shaping the lightcurves generate X-ray radiation that could be followed with XMM-Newton (Jansen et al., 2001) and SWIFT Gehrels et al. 2004.
@@
Including the effect of jets in simulating the in-spiral of the compact object inside the giant and modeling their interaction with the ejecta and CSM might yield different results for the lightcurves. Modeled jet-shaped lightcurves exhibit higher luminosity peaks compared to a spherical explosion as a result of the extra energy source and relativistic beaming, and more abrupt declines due to a faster polar ejecta (e.g., Kaplan & Soker 2020a). The aspherical nature of jet launching would lead to a highly asymmetrical CSM, variability in the optical and X-ray emissions and a bumpy lightcurve (see Schreier et al. 2021 for lightcurves powered by jets that are launched when an NS crosses the envelope of a giant star in an eccentric orbit). The negative jet feedback mechanism (e.g., Grichener et al. 2021; Hillel et al. 2022b; see section 3) would lead to changes in the accretion rate by the compact object and therefore in the jets luminosity, contributing to the variability in the emission. The large amount of mass expelled due to jet launching during the merger would expand faster than in a case with no jets, and could lead to enhanced dust formation with strong IR radiation that could be observed with JWST (Gardner et al., 2006) and the upcoming Nancy Grace Roman Telescope. In cases where the merger occurs after the ejection of most of the envelope, the jets could stream out freely and give rise to a LGRB in the wavelength range of SWIFT(Gehrels et al. 2004; see section 5.1.5). In this case the ejecta would exhibit low amounts of hydrogen, as in Type Ibn (e.g., Pastorello et al. 2007) and Icn SN (e.g., Gal-Yam et al. 2021).Further study and modelling are needed to determine the role jets would have in shaping the lightcurves of NS/BH-core mergers, and give prediction of observational signatures to search for them in existent data and future surveys.
@@ 5.1.2 Core collapse supernovae
Barkov & Komissarov (2011) ties NS-core mergers to hypernova explosions. They propose that the high mass accretion rates during the merger can lead to the generation of strong magnetic fields through differential rotation energizing a very luminous Type II SN with a high plateau followed by a steep decline. The hydrogen features come from the giant’s envelope, and the sharp drop in the brightness is due to low amounts of synthesized is such an explosion that occur in relatively low densities and the fast spin-down of the NS. We note that the spinning of the NS in this scenario, and hence its potential to produce a magnetically driven explosion during the merger, is very sensitive to the specific angular momentum of the matter accreted by the NS itself. If the vast accretion leads the NS to collapse into a BH before the magnetically driven explosion manages to occur, then the relativistic jets launched due to the accretion by the BH will most likely remain chocked in the inflated envelope of the companion (e.g., Grichener & Soker 2021), leading to a Type II SN.
@@ 5.1.3 Thermonuclear explosion
In cases where the tidal disruption radius of the core is outside of the core (third row of table 1), the NS/BH will exert tidal forces that disrupt the core before it reaches the core’s surface. In this case, the accretion disk formed around the compact object could experience thermonuclear outbursts, powering a bright transient (e.g., Ginat et al. 2020; Grichener & Soker 2023), with some similarities to the disruption of a white dwarf by a compact object (e.g., Fryer et al. 1999; Zenati et al. 2019; Zenati et al. 2020). Grichener & Soker 2023 propose that this kind of NS-core thermonuclear merger could be the origin of the enigmatic SN remnant W49B. The ignition of thermonuclear burning in the non-axisymmetric disk could account for the highly uneven distribution of metals, and its peculiar morphology (see Fig 5) could be attributed to jets-ejecta interactions (Akashi et al., 2018). If the nuclear burning and jet-launching occurs relatively early in the tidal disruption driven merger, the disk will not be fully symmetric and the disrupted material will be concentrated in two spiral tails, leading to its ejection in an elongated structure.
@@ 5.1.4 Luminous fast blue optical transients
Several studies suggested that LFBOTs could originate in the mergers of NSs or BHs with helium cores of giant stars (Soker et al. 2019; Soker 2022; Metzger 2022; Cohen & Soker 2023; Tuna & Metzger 2023). While the classical prompt NS/BH-core merger scenario during (Soker et al., 2019) or shortly after (Soker, 2022) CEE can explain both the presence of a powering compact object and a massive CSM around the LFBOT, it has a problem accounting for the lack of hydrogen in the CSM, as hydrogen-rich debris of the common envelope would probably give rise to hydrogen features in the transient spectra. Metzger (2022) suggest that LFBOTs originate in the merger of an NS or a BH with the naked core of the giant. 555If sufficiently massive and luminous, the naked core might appear as an hydrogen-poor classical Wolf-Rayet star. See Shenar 2024 for a recent review. The NS/BH-core binary system could be formed following the ejection of the common envelope at the end of the CEE phase (stage 7 in channel I, left panel of Fig. 1) or following a stable mass transfer episode that leads to Roche lobe Overflow (RLOF) and could also power fast radio bursts at an early stage (e.g., Sridhar et al. 2021). The merger is delayed by about yr compared to the mass transfer event and occurs due to the interaction of the surviving binary system with a circumbinary disk that is formed around the system. Loss of angular momentum to the disk through viscous torques causes the orbit of the binary to shrink until the compact object merges with the core when most of the envelope matter is far from the location of the merger, explaining the small amounts of hydrogen in the CSM. The dense CSM can come from mass that the core loses at the beginning of the merger and from winds due to the photo-evaporation of the circumbinary disk. The disk-wind outflow launched during the merger could explain both the asymmetry of the ejecta with the different velocity components that come from the non-spherical nature of these outflows, and the low amounts of synthesized in the inner regions of the disk where the accretion rates and temperatures are high enough. Since the compact object merges with the helium core of the giant, we do not expect self-neutronization of the disk and r-process nucleosynthesis in these events.
@@
Fig. 6 (taken from Metzger 2022) shows a good comparability between the total optical (solid black line) and X-ray (solid dark blue line) lightcurves powered by a naked core-BH delayed merger model of an equal mass binary of and the observed AT2018cow energies from Margutti et al. (2019) in the respective bands (black, dark blue and light blue dots for optical, soft X-ray and hard X-ray, respectively). The optical luminosity is powered by both the shock interaction between the ejecta and CSM (black dashed line) and re-processing of the X-ray emission from the disk (black dotted-dashed line). The luminosity of the X-rays that manage to escape and avoid being re-processed is presented by the solid dark blue line, that might flatten (dotted blue line) as the BH mass accretion rate reaches the Eddington value (the Eddington luminosity is presented by the brown dotted-dashed line) as seen in X-ray binaries. While the steepening of the X-ray light curve around a month is not obtained by this toy-model, it could be potentially explained by a reduced mass accretion rate onto the BH due to expansion of the disk.

### v2

@@ 4 R-process nucleosynthesis and enrichment
The possible formation site(s) of r-process elements remain a highly debated subject (e.g., Hotokezaka et al. 2018; Siegel et al. 2019; Ji et al. 2019; Banerjee et al. 2020; Chen et al. 2024a; Chen et al. 2024b; Tsujimoto 2024). In 2017, LIGO detected the first GW signature from binary NS (BNS) mergers, commonly known as GW170817 (Abbott et al., 2017b). The electromagnetic counterpart that was associated with this GW event (kilonova AT2017gfo; Abbott et al. 2017a) showed traces of lanthanides in their equatorial outflow (Kasen et al., 2017), hence confirming r-process nucleosynthesis in BNS mergers. Despite this groundbreaking progress, which was further strengthened by an additional discovery of a kilonova that might be associated with another BNS merger with JWST (Levan et al., 2024), it is not clear whether BNS mergers are the only site of heavy element formation in the Universe. The long delay times until the merger and subsequent r-process nucleosynthesis, reaching tens of Gyrs (e.g., Belczynski et al. 2018), might imply there are not enough mergers to account for heavy element formation in the early Universe (e.g., Côté et al. 2019; Kobayashi et al. 2023. See Maoz & Nakar 2024 and Beniamini & Piran 2024 for a different view). Moreover, these long delay times combined with the two natal-kicks that the binary system receives following each SN explosion (see left panel of Fig. 1) might drive the BNS outside of ultra-faint dwarf (UFD) galaxies before they manage to merge, implying that they are not able to enrich them with heavy elements (e.g., Bonetti et al. 2019). However, recent observations of the UFD galaxy Reticulum II show r-process enhanced gas (e.g., Ji et al. 2016; Roederer et al. 2016; Ji et al. 2023; Simon et al. 2023), highlighting the need for additional r-process nucleosynthesis sites (see Beniamini et al. 2016 for a another view.)
@@
Papish et al. (2015) suggest that r-process nucleosynthesis can occur during the merger of an NS with the core of a giant star. As the NS spirals through the core of the giant, it might accrete mass at very high rates (see section 3 for a discussion about the accretion rate). The vast accretion destroys the core forming a thick accretion disk around the NS that can self neutronize through electron captures due to the extremely high temperatures and densities in the disk, providing a neutron-rich environment where heavy elements can be formed via rapid neutron captures by isotopes that come from outer cooler layers of the disk. Grichener & Soker (2019) showed that under their model assumptions the matter of the accretion disk could be neutron-rich enough for r-process nucleosynthesis if the merger occurs after helium core depletion, when the core is CO-rich. They perform one-dimensional stellar evolution simulations of several massive stars that engulf an NS in different stages of their evolution and mimic the effect of the NS on the giant by removing envelope mass from the location of the NS outwards. Using the Bondi-Hoyle-Lyttleton approximation (see equation 2), they compute the accretion rate of the NS as it merges with the core of the giant, and then apply it to estimate the typical temperatures and densities of the material in the disk within the analytical framework of a neutrino cooled accretion disk from Chevalier (1996)
@@
Since the orbital shrinkage leading to mergers of NSs with giants’ cores is driven by dynamical friction during CEE rather than GW irradiation, they occur on a dynamical timescale and have much shorter delay time from the formation of the binary compared to BNS mergers ( tens of Myrs; from the population synthesis data of Grichener 2023; comparable to the lifetime of the NS progenitor). These shorter delay times offer an advantage in explaining the presence of heavy element in the early Universe. Moreover, NS-core mergers suffer from a single natal-kick induced by the SN that formed the NS, increasing the possibility that they will be retained and produce heavy elements in UFD galaxies (Grichener & Soker, 2022). Grichener (2023) computes the event rates of NS-core mergers with CO-rich cores and find it to be around 1 event per 1000 CCSNe for fiducial model parameters, which is comparable to the rate of BNS mergers inferred by LIGO (Abbott et al., 2023), and sufficient to account for a substantial fraction of observed r-process abundances. Grichener et al. (2022) use a Galactic Chemical Evolution model of the Milky-Way (Kobayashi et al., 2020) and include the contribution of NS-core mergers to r-process nucleosynthesis. They study the time evolution of europium in the Galaxy and find that europium nucleosynthesis yields in NS-core mergers (orange dots in Fig. 3) are within the confines of the observed values at earlier times and they can qualitatively reproduces the ’evolution knee’ that follows, while BNS mergers (yellow dots in Fig. 3) struggle to do so.
@@ 5.1.1 Merger-driven lightcurves properties
While three-dimensional hydrodynamical simulations that follow the the outflows and outgoing radiation during the merger of the compact object with the core are scarce, simulating the final CEE stages of a compact object and a giant star has value in building intuition. Schrøder et al. (2020) model the in-spiral of a BH inside a giant and calculate the lightcurves expected prior to the merger of the BH with the core accounting for the effects of CSM interaction with a spherically generated shock instead of jets. They continue the simulations until the orbital separation of the binary is about , assuming the subsequent merger is expected to occur rapidly on a timescale of less then a year. They then analyze the ejection of matter during the final stage of CEE and find the resultant CSM distribution . We note that by not modelling the accretion onto the BH, their resultant CSM might be more massive than expected in such a merger. For computational simplicity, they derive a spherically symmetric density distribution of the ejecta and model the CSM interactions using a one-dimensional radiative hydrodynamics code.
@@
As can be seen in Fig 4, Schrøder et al. (2020) find that at early times the lightcurves are mostly dominated by CSM interactions. As in other models of dense CSM close to the giant star, they find fast rising high luminosities that reach plateau over a period of hundreds of days. 333In the case of an extended CSM the decay of the lightcurve can happen over a period of thousands of days (e.g., Smith 2017). Typical models have higher luminosities than type IIp SNe, but still lower than in the case of super-luminous SNe. The top panel of Fig. 4 shows that more energetic mergers result in shorter, more luminous transients with less prominent effect of CSM interactions on the lightcurve as the larger photospheres of the explosive transients are further out from the CSM edge. The lightcurve of the most energetic transient ( ergs) resembles the shape of a type IIp SN (Schrøder et al., 2020). Higher giant masses (middle panel of Fig. 4) result in lower ejecta velocities that lead to longer transients. Different giant radii (bottom panel of Fig. 4) affect the transients mainly by forming CSMs of different extents. Smaller giants lead to more compact CSMs that can affect the lightcurve through interactions at earlier times. Such explosive transients could be detected by current facilities, such as ZTF (Bellm et al., 2019) and ASAS-SN (Kochanek et al., 2017), as well as with future programs like LSST. The CSM interactions involved in shaping the lightcurves generate X-ray radiation that could be followed with XMM-Newton (Jansen et al., 2001) and SWIFT Gehrels et al. 2004.
@@
Including the effect of jets in simulating the in-spiral of the compact object inside the giant and modeling their interaction with the ejecta and CSM might yield different results for the lightcurves. Modeled jet-shaped lightcurves exhibit higher luminosity peaks compared to a spherical explosion as a result of the extra energy source and relativistic beaming, and more abrupt declines due to a faster polar ejecta (e.g., Kaplan & Soker 2020a). The aspherical nature of jet launching would lead to a highly asymmetrical CSM, variability in the optical and X-ray emissions and a bumpy lightcurve (see Schreier et al. 2021 for lightcurves powered by jets that are launched when an NS crosses the envelope of a giant star in an eccentric orbit). The negative jet feedback mechanism (e.g., Grichener et al. 2021; Hillel et al. 2022b; see section 3) would lead to changes in the accretion rate by the compact object and therefore in the jet luminosity, contributing to the variability in the emission. The large amount of mass expelled due to jet launching during the merger would expand faster than in a case with no jets, and could lead to enhanced dust formation with strong IR radiation that could be observed with JWST (Gardner et al., 2006) and the upcoming Nancy Grace Roman Telescope. In cases where the merger occurs after the ejection of most of the envelope, the jets could stream out freely and give rise to a LGRB in the wavelength range of SWIFT(Gehrels et al. 2004; see section 5.1.5). In this case the ejecta would exhibit low amounts of hydrogen, as in Type Ibn (e.g., Pastorello et al. 2007) and Icn SN (e.g., Gal-Yam et al. 2021).Further study and modelling are needed to determine the role jets would have in shaping the lightcurves of NS/BH-core mergers, and give prediction of observational signatures to search for them in existent data and future surveys.
@@ 5.1.2 Core collapse supernovae
Barkov & Komissarov (2011) ties NS-core mergers to hypernova explosions. They propose that the high mass accretion rates during the merger can lead to the generation of strong magnetic fields through differential rotation energizing a very luminous Type II SN with a high plateau followed by a steep decline. The hydrogen features come from the giant’s envelope, and the sharp drop in the brightness is due to low amounts of synthesized is such an explosion that occur in relatively low densities and the fast spin-down of the NS. We note that the spinning of the NS in this scenario, and hence its potential to produce a magnetically driven explosion during the merger, is very sensitive to the specific angular momentum of the matter accreted by the NS itself. If the vast accretion leads the NS to collapse into a BH before the magnetically driven explosion manages to occur, then the relativistic jets launched due to the accretion by the BH will most likely remain choked in the inflated envelope of the companion (e.g., Grichener & Soker 2021), leading to a superluminous Type II SN. Ablimit et al. 2022 suggest that a similar transient can arise from the collapse of a white-dwarf to a NS or a BH as it mergers with the core of a giant star during CEE.
@@ 5.1.3 Thermonuclear explosion
In cases where the tidal disruption radius of the core is outside of the core (third row of table 1), the NS/BH will exert tidal forces that disrupt the core before it reaches the core’s surface. In this case, the accretion disk formed around the compact object could experience thermonuclear outbursts, powering a bright transient (e.g., Ginat et al. 2020; Grichener & Soker 2023), with some similarities to the disruption of a white dwarf by a compact object (e.g., Fryer et al. 1999; Zenati et al. 2019; Zenati et al. 2020). Grichener & Soker 2023 propose that this kind of NS-core thermonuclear merger could be the origin of the enigmatic SN remnant W49B. The ignition of thermonuclear burning in the non-axisymmetric disk could account for the highly uneven distribution of metals, and its peculiar morphology (see Fig 5) could be attributed to jet-ejecta interactions (Akashi et al., 2018). If the nuclear burning and jet-launching occurs relatively early in the tidal disruption driven merger, the disk will not be fully symmetric and the disrupted material will be concentrated in two spiral tails, leading to its ejection in an elongated structure.
@@ 5.1.4 Luminous fast blue optical transients
Several studies suggested that LFBOTs could originate in the mergers of NSs or BHs with helium cores of giant stars (Soker et al. 2019; Soker 2022; Metzger 2022; Cohen & Soker 2023; Tuna & Metzger 2023). While the classical prompt NS/BH-core merger scenario during (Soker et al., 2019) or shortly after (Soker, 2022) CEE can explain both the presence of a powering compact object and a massive CSM around the LFBOT, it has a problem accounting for the lack of hydrogen in the CSM, as hydrogen-rich debris of the common envelope would probably give rise to hydrogen features in the transient spectra. Metzger (2022) suggest that LFBOTs originate in the merger of an NS or a BH with the naked core of the giant. 555If sufficiently massive and luminous, the naked core might appear as an hydrogen-poor classical Wolf-Rayet star. See Shenar 2024 for a recent review. The NS/BH-core binary system could be formed following the ejection of the common envelope at the end of the CEE phase (stage 7 in channel I, left panel of Fig. 1) or following a stable mass transfer episode that leads to Roche lobe Overflow (RLOF) and could also power fast radio bursts at an early stage (e.g., Sridhar et al. 2021). The merger is delayed by about yr compared to the mass transfer event and occurs due to the interaction of the surviving binary system with a circumbinary disk that is formed around the system (for a population synthesis study that explores the conditions for the formation of circumbinary disks in post common envelope binary systems with compact objects see Unger et al. 2024). Loss of angular momentum to the disk through viscous torques causes the orbit of the binary to shrink until the compact object merges with the core when most of the envelope matter is far from the location of the merger, explaining the small amounts of hydrogen in the CSM. The dense CSM can come from mass that the core loses at the beginning of the merger and from winds due to the photo-evaporation of the circumbinary disk. The disk-wind outflow launched during the merger could explain both the asymmetry of the ejecta with the different velocity components that come from the non-spherical nature of these outflows, and the low amounts of synthesized in the inner regions of the disk where the accretion rates and temperatures are high enough. Since the compact object merges with the helium core of the giant, we do not expect self-neutronization of the disk and r-process nucleosynthesis in these events.
@@
Fig. 6 (taken from Metzger 2022) shows a good compatibility between the total optical (solid black line) and X-ray (solid dark blue line) lightcurves powered by a naked core-BH delayed merger model of an equal mass binary of and the observed AT2018cow energies from Margutti et al. (2019) in the respective bands (black, dark blue and light blue dots for optical, soft X-ray and hard X-ray, respectively). The optical luminosity is powered by both the shock interaction between the ejecta and CSM (black dashed line) and re-processing of the X-ray emission from the disk (black dotted-dashed line). The luminosity of the X-rays that manage to escape and avoid being re-processed is presented by the solid dark blue line, that might flatten (dotted blue line) as the BH mass accretion rate reaches the Eddington value (the Eddington luminosity is presented by the brown dotted-dashed line) as seen in X-ray binaries. While the steepening of the X-ray light curve around a month is not obtained by this toy-model, it could be potentially explained by a reduced mass accretion rate onto the BH due to expansion of the disk.